Evaluate the deteminant
1 0 -1
-2 3 5
6 -3 4

Answer by fractalier(6550)

1 0 -1
-2 3 5
6 -3 4
we have
1(3*4 - (-3)*(5) + 0 + (-1)((-2)(-3) - (3)(6)) =
(12 + 15) - (6 - 18) =
27 + 12 =
39